# MoE-Loco: Mixture of Experts for Multitask Locomotion

### v1

@@ III-B MoE Based Multitask Locomotion Learning
The MoE architecture facilitates the coordination of similar task skills while minimizing conflicts between heterogeneous tasks by dynamically routing tasks to appropriate experts. This automatic routing enables specialization, improving both efficiency and task performance. Additionally, we incorporate MoE into the critic network to better capture diverse task reward structures. The actor MoE, critic MoE, and gating network share the same input $h_{t}$, which encodes proprioceptive states and task-specific features. By using a shared gating network, we ensure consistency between policy evaluation and action generation.
@@ V-C Network Architecture Details
Network Type Hidden dims

### v2

@@ III-B MoE Based Multitask Locomotion Learning
The MoE architecture facilitates the coordination of similar task skills while minimizing conflicts between heterogeneous tasks by dynamically routing tasks to appropriate experts. This automatic routing enables specialization, improving both efficiency and task performance. Additionally, we incorporate MoE into the critic network to better capture diverse task reward structures. The actor MoE and critic MoE shares the same gating network. By using a shared gating network, we ensure consistency between policy evaluation and action generation.
@@ V-C Network Architecture Details
Network Type Dims
@@
V-D Domain Randomization
We introduce dynamic randomization to ensure our policy can be safely transfered to real environment. The random ranges are shown in the Table VII.
TABLE VII: Domain randomization
Parameters Range Unit
Base mass [1, 3] $kg$
Mass position of X axis [-0.2, 0.2] $m$
Mass position of Y axis [-0.1, 0.1] $m$
Mass position of Z axis [-0.05, 0.05] $m$
Friction [0, 2] -
[…]
Proprioception latency [0.005, 0.045] $s$
Gaussian noise are added to the input observation to simulate the noise in the real world, as shown in  Table VIII.
TABLE VIII: Gaussian noise
Observation Gaussian Noise Amplitude Unit
Linear velocity 0.05 $m/s$
Angular velocity 0.2 $rad/s$
Gravity 0.05 $m/s^{2}$
Joint position 0.01 $rad$
Joint velocity 1.5 $rad/s$
V-E Training Details
In order to prevent legs dragging on the floor, and achieve a robust performance on uneven terrains, we apply fractal noise [7] to the ground. We use maximum noise scale $\text{z}_{max}=0.1$. We use PPO as our reinforcement learning algorithm. The hyperparameters are shown in Table IX.
TABLE IX: PPO hyperparameters
      Hyperparameter       Value
      clip min std       0.05
      clip param       0.2
      gamma       0.99
      lam       0.95
      desired kl       0.01
      entropy coef       0.01
      learning rate       0.001
      max grad norm       1
      num mini batch       4
      num steps per env       24